# Data-Efficient Extremum-Seeking Control Using Kernel-Based Function Approximation

### v1

@@ 5 Simulation study
with $\theta^{*}=[\begin{array}[]{@{}cc@{}}\theta_{1}^{*}&\theta_{2}^{*}\end{array}]^{\intercal}$. For each standard update step (in either approach), we use a gradient-based update step, where the gradient of the input-output map is estimated using the central difference scheme. To this end, we perform $n_{v}=4$ experiments per update step, during which constant dither $v_{j}$, $j=1,\dots,n_{v}$, with amplitude $c_{v}$ is added to the current optimizer state $\widehat{\theta}_{k}$ along the positive and negative axes of the parameters $\theta_{1}$ and $\theta_{2}$, respectively, to obtain the inputs $\widetilde{\theta}_{N_{k}+j}$ as in (14), i.e., $v_{1}$ to $v_{4}$ are given as
@@
The corresponding output measurements $\widetilde{y}_{N_{k}+j}$ (cf. (15)), are then used in an update step of the form
@@
$\Gamma$ $c$ $\underline{\mu}$ $\overline{\mu}$ $\varrho$
@@
Starting both approaches with initial conditions $x(0)=[\begin{array}[]{@{}cc@{}}0&0\end{array}]^{\intercal}$ and $\widehat{\theta}_{0}=[\begin{array}[]{@{}cc@{}}-2&-4\end{array}]^{\intercal}$, shows that using the kernel-based approach the optimizer state $\widehat{\theta}$ converges with 60 experiments to the same neighborhood of the minimizer $\theta^{*}$ as the standard approach reaches with 100 experiments, as illustrated by Figure 2. The number of experiments needed to reach a small neighborhood of the minimizer is thus reduced by 40%. This reduction is the result of the fact that three of the update steps performed in the kernel-based approach ($k=3,5$, and $7$) are kernel-based update steps, as illustrated in Figure 3, which do not require additional experiments to be performed to update the parameters and which allow larger steps to be taken due to the larger computed optimizer gain ($\mu=14.1,26.6$, and $17.4$, respectively) than the fixed value used in the standard update steps ($\widetilde{\mu}=5$). As a result of the larger optimizer gain, the optimizer state $\widehat{\theta}$ also converges in fewer update steps to the same neighborhood of the minimizer $\theta^{*}$ than the standard approach: 18 instead of 25 updates, which is a reduction of 28%.
@@
Performing these large kernel-based steps is possible because the approximation $m_{k}(\theta)$ quickly becomes an accurate representation of the true input-output map $f(\theta)$ along the search direction. To illustrate this, we note that most of the inputs $\widetilde{\theta}_{i}$ applied during the experiments in both approaches lie close to the line through the initial optimizer state $\widehat{\theta}_{0}$ and the minimizer $\theta^{*}$, as shown in Figure 4. Plotting the cross-section of the input-output map $f(\theta)$ and the approximation $m_{k}(\theta)$ at the three kernel-based update steps ($k=3,5$ and $7$) along this line indeed shows that along this search direction the approximation $m_{k}(\theta)$ quickly becomes an accurate approximation of $f(\theta)$, as shown in Figure 5.
@@
Figure 2: The optimizer state $\widehat{\theta}$ for both ESC approaches as a function of the number of experiments. The proposed kernel-based approach requires only 60 experiments to reach the same neighborhood of the optimum that the standard approach reaches with 100 experiments.
@@
Figure 3: The optimizer state $\widehat{\theta}$ for both ESC approaches as a function of the number of update steps. The third, fifth, and seventh update in the proposed kernel-based approach are kernel-based update steps, which do not require additional experiments to be performed to perform the update and allow searching for a suitable optimizer gain, resulting in needing only 18 update steps to reach the same neighborhood of the optimum as the standard approach does with 25 update steps.
@@
Figure 4: Contour plot of the steady-state input-output map $f(\theta)$, along with inputs $\widetilde{\theta}_{i}$ applied during experiments in both approaches. These applied inputs lie close to the line passing through the initial optimizer state $\widehat{\theta}_{0}$ and the minimizer $\theta^{*}$ (shown in green).
@@
Figure 5: Cross-section of the input-output map $f(\theta)$ as well as the approximation $m_{k}(\theta)$ at different kernel-based update steps, along the search direction (cf. green line in Figure 4). $m_{k}(\theta)$ quickly becomes an accurate approximation of $f(\theta)$ allowing large kernel-based update steps to be taken from the circle to the square.
@@ 6 Conclusion
Motivated by the typically slow convergence in extremum-seeking control and the potentially costly nature of performing experiments in practical applications, we presented a novel extremum-seeking approach aimed at increasing the convergence rate and reducing the total number of experiments needed to optimize performance. The proposed approach uses kernel-based function approximation to construct online an approximation of the steady-state input-output map of the system, based on data collected during regular extremum-seeking steps. This approximation is used to determine a suitable search direction and optimizer gain to perform a parameter update without performing additional experiments whenever it is sufficiently accurate to guarantee a decrease in the input-output map. If such decrease cannot be guaranteed, a regular extremum-seeking parameter update step is performed to update the parameters, and the data collected during this update step is used to improve the approximation for the next parameter update step. By using this approach, reductions in both the number of performed experiments and the number of update steps can be obtained, as illustrated by a simulation study in which the number of experiments and updates are reduced by 40% and 28%, respectively.
Future research could consider extensions to a more general class of optimization methods as in Khong et al. (2013a, b), for example, for both the standard parameter update steps and kernel-based update steps. Other potential research directions include the use of variable waiting times instead of a fixed one, as, e.g., in Poveda and Teel (2017), to speed up convergence even further, or extensions in which only a subset of all collected data are used, e.g., to reduce computational complexity or to deal with input-output maps that are slowly changing over time.
@@ Appendix A Proof of Theorem 10
We start the proof by showing that under the conditions in Theorem 10, the bound $\overline{b}(\theta,\mu)$ in (39) is indeed the maximum value that $f(\theta^{+})-f(\theta)+c\mu\nabla f(\theta)^{\intercal}\nabla m_{k}(\theta)$ can obtain given the input $\theta\in\mathbb{R}^{n_{\theta}}$. That is, we will start by showing that (39) is the same as
$\overline{b}(\theta,\mu)=$
@@
:=$\mathcal{H}^{\parallel}:=\{f\in\;$ $\mathcal{H}:f\in{\rm span}\{\kappa(\cdot,\widetilde{\theta}_{1}),\dots,\kappa(\cdot,\widetilde{\theta}_{N_{k}}),\kappa(\cdot,\theta),$
@@
Let :=$\mathcal{H}^{\perp}:=\{g\in\mathcal{H}:\langle g,f^{\parallel}\rangle=0,\,\forall f^{\parallel}\in\mathcal{H}^{\parallel}\}$ be the subspace of functions in $\mathcal{H}$ that are orthogonal to $\mathcal{H}^{\parallel}$ in (51). It then follows that $\mathcal{H}=\mathcal{H}^{\parallel}\oplus\mathcal{H}^{\perp}$, with $\oplus$ denoting the vector direct sum, and that for all $f\in\mathcal{H}$ there exist $f^{\parallel}\in\mathcal{H}^{\parallel}$ and $f^{\perp}\in\mathcal{H}^{\perp}$ such that $f=f^{\parallel}+f^{\perp}$. Using this decomposition of $f$, and observing that $\langle f^{\perp},\kappa(\cdot,z)\rangle$ and $\langle f^{\perp},D^{(0,e_{j})}\kappa(\cdot,\theta)\rangle$ are zero for $z\in\{\widetilde{\theta}_{1},\dots,\widetilde{\theta}_{N_{k}},\theta,\theta^{+}\}$ and $j=1,\dots,n_{\theta}$, respectively, since for those $z$ and $j$ it holds that $\kappa(\cdot,z)\in\mathcal{H}^{\parallel}$, and $D^{{(0,e_{j})}}\kappa(\cdot,\theta)\in\mathcal{H}^{\parallel}$, we obtain using (52) and (53) that (50a) can be written as
@@
$\>\lvert f^{\parallel}(\widetilde{\theta}_{i})-\widetilde{y}_{i}\rvert\leq\bar{\delta}\quad\forall i\in\{1,\dots,N_{k}\}.$ (57c)
@@
Step 2: Next, we address the attainment of the supremum by showing that the constraints in (57) define a closed and bounded feasible set. First, note that $\lVert f^{\parallel}\rVert^{2}\leq\Gamma^{2}$ defines a closed and bounded set, since it is the sublevel set of a norm. Next, we note that since sets of the form $\{a\in\mathbb{R}:\lvert a\rvert\leq b\}$ are closed in $\mathbb{R}$, so are $\{f^{\parallel}(\widetilde{\theta}_{i})\in\mathbb{R}:\lvert f^{\parallel}(\widetilde{\theta}_{i})-\widetilde{y}_{i}\rvert\leq\bar{\delta}\}$ $\forall i\in\{1,\dots,N_{k}\}$. Furthermore, the evaluation functional $L_{z}(f^{\parallel})=f^{\parallel}(z)$ is a linear operator and thus the pre-images $L_{z}^{-1}$ of these closed sets are also closed. Consequentially, $\{f^{\parallel}\in\mathcal{H}^{\parallel}:\lvert f^{\parallel}(\widetilde{\theta}_{i})-\widetilde{y}_{i}\rvert\leq\bar{\delta}\}$ $\forall i\in\{1,\dots,N_{k}\}$ are closed in $\mathcal{H}^{\parallel}$. Since the intersection of a finite number of closed sets is necessarily closed, the constraints (57b)-(57c) define a closed feasible set. Moreover, this set is bounded because it is contained in the bounded set defined by the constraint (57b). Since $\mathcal{H}^{\parallel}$ in (51) is finite dimensional, any closed and bounded subset is compact by the Heine-Borel theorem (see, e.g., Royden and Fitzpatrick (2010, Theorem 20)). Therefore, by the extreme value theorem, the continuous objective (57a) attains a maximum on the feasible compact set. Moreover, from Step 1 it follows that the optimizer for which this maximum is obtained must be in $\mathcal{H}^{\parallel}$ ($f^{\perp}=0$), whose members $f^{\parallel}\in\mathcal{H}^{\parallel}$ by (51) have the form
@@
with $K_{\Omega}(\cdot)$ as in (29) for the inputs $\Omega=\{\widetilde{\theta}_{1},\dots,\widetilde{\theta}_{N_{k}}\}$, $\nabla K(\theta,\cdot)$ as in (37), and $\beta\in\mathbb{R}^{N_{k}+2+n_{\theta}}$ a weight vector.
@@
$\begin{bmatrix}f^{\parallel}(\widetilde{\theta}_{1})&\cdots&f^{\parallel}(\widetilde{\theta}_{N_{k}})&f^{\parallel}(\theta)&f^{\parallel}(\theta^{+})&\nabla f(\theta)^{\intercal}\end{bmatrix}^{\intercal}$
@@
for all $i\in\{1,\dots,N_{k}\}$ with $\xi_{i}\in\mathbb{R}^{N_{k}+2+n_{\theta}}$ a vector with $i$-th element equal to one and other elements equal to zero.
Using (60)-(62), and the observation from Step 2 that the objective (57a) attains a maximum on the feasible compact set, we obtain (39) from (57), completing the proof for $\overline{b}(\theta,\mu)$.
The proof that under the conditions in Theorem 10 $\underline{b}(\theta)$ in (38) is a lower bound on $\nabla f(\theta)^{\intercal}\nabla m_{k}(\theta)$, i.e., that (38) is the same as
$\underline{b}(\theta)=$
@@
$\lvert f(\widetilde{\theta}_{i})-\widetilde{y}_{i}\rvert\leq\bar{\delta}\quad\forall i\in\{1,\dots,N_{k}\},$ (63c)
@@ Appendix B Proof of Lemma 13
The proof of Lemma 13 is as follows. Since by the conditions of the lemma $\overline{b}(\theta,\mu)\leq 0$, it follows from Theorem 10 that the inequality $f(\theta^{+})\leq f(\theta)-c\mu\nabla f(\theta)^{\intercal}\nabla m_{k}(\theta)$ with $c\in(0,1)$ is satisfied. Therefore, using the definitions of $V(\cdot)$ and $\Delta V(\cdot)$ in (40) and (44), we obtain that the inequality
@@
holds. Furthermore, since by the conditions of the lemma $\underline{b}(\theta)>0$ for all $\𝒞\theta\in\mathbb{R}\backslash\mathcal{C}$, and by Assumption 2(iii) $\nabla f(\theta)=0$ if and only if $\theta\in\mathcal{C}$, it follows from Theorem 10 and (64) that $\Delta V(\theta)<0$ for all $\theta\in\mathbb{R}^{n_{\theta}}\backslash\mathcal{C}$ and $\Delta V(\theta)=0$ for all $\theta\in\mathcal{C}$. Hence, for any $\Delta_{\theta}$ such that $\lVert\theta\rVert_{\mathcal{C}}\leq\Delta_{\theta}$, there exists a class-$\mathcal{K}$ function $\bar{\rho}_{k}$ such that $\bar{\rho}_{k}(\lVert\theta\rVert_{\mathcal{C}})\leq-\Delta V(\theta)$ (see, e.g., Khalil (2002, Lemma 4.3)). Moreover, it follows from the definition of $V(\cdot)$ in (40) and Assumption 10 that $\lVert\theta\rVert_{\mathcal{C}}\geq\omega_{2}^{-1}(V(\theta))$ and thus that $-\bar{\rho}_{k}(\lVert\theta\rVert_{\mathcal{C}})\leq-\bar{\rho}_{k}(\omega_{2}^{-1}(V(\theta)))$. Therefore, we obtain from (64) that during kernel-based extremum-seeking steps the inequality

### v2

@@ 5 Simulation study
with $\theta^{*}=[\begin{array}[]{@{}cc@{}}\theta_{1}^{*}&\theta_{2}^{*}\end{array}]^{\intercal}$. For each standard update step (in either approach), we use a gradient-based update step, where the gradient of the input-output map is estimated using the central difference scheme. To this end, we perform $n_{v}=4$ measurements per update step, during which constant dither $v_{j}$, $j=1,2,\dots,n_{v}$, with amplitude $c_{v}$ is added to the current optimizer state $\widehat{\theta}_{k}$ along the positive and negative axes of the parameters $\theta_{1}$ and $\theta_{2}$, respectively, to obtain the inputs $\widetilde{\theta}_{N_{k}+j}$ as in (10), i.e., $v_{1}$ to $v_{4}$ are given as
@@
We use the corresponding output measurements $\widetilde{y}_{N_{k}+j}$ (cf. (11)), in an update step of the form
@@
$\Gamma$ $c$ $\underaccent{\bar}{\mu}$ $\bar{\mu}$ $\varrho$
@@
Starting both approaches with initial conditions $x(0)=[\begin{array}[]{@{}cc@{}}0&0\end{array}]^{\intercal}$ and $\widehat{\theta}_{0}=[\begin{array}[]{@{}cc@{}}-2&-4\end{array}]^{\intercal}$, shows that using the kernel-based approach the optimizer state $\widehat{\theta}$ converges with 60 measurements to the same neighborhood of the minimizer $\theta^{*}$ as the standard approach reaches with 100 measurements, as illustrated by Figure 2. The number of measurements needed to reach a small neighborhood of the minimizer is thus reduced by 40%. This reduction is the result of two benefits that the kernel-based approach provides. First, three of the update steps performed in the kernel-based approach ($k=3,5$, and $7$) are kernel-based update steps, as illustrated in Figure 3, which do not require additional measurements to be performed to update the parameters. Second, these kernel-based update steps are larger due to the use of a larger optimizer gain ($\mu=14.1$, $26.6$, and $17.4$, respectively) computed by the backtracking line search, compared to the fixed value used in the standard update steps ($\widetilde{\mu}=5$). Because of these larger steps, the optimizer state $\widehat{\theta}$ converges in fewer update steps to the same neighborhood of the minimizer $\theta^{*}$ than the standard approach: 18 instead of 25 updates, which is a reduction of 28%.
@@
Performing these large kernel-based steps is possible because by using the input-output data collected during regular operation of the extremum-seeking controller to construct and update the approximation $m_{k}(\theta)$ online, it quickly becomes an accurate representation of the true input-output map $f(\theta)$ along the search direction. To illustrate this, we note that most of the inputs $\widetilde{\theta}_{i}$ applied during the measurements in both approaches lie close to the line through the initial optimizer state $\widehat{\theta}_{0}$ and the minimizer $\theta^{*}$, as shown in Figure 4. Plotting the cross-section of the input-output map $f(\theta)$ and the approximation $m_{k}(\theta)$ at the three kernel-based update steps ($k=3,5$ and $7$) along this line indeed shows that along this search direction the approximation $m_{k}(\theta)$ quickly improves, as shown in Figure 5.
@@
Figure 2: The optimizer state $\widehat{\theta}$ for both ESC approaches as a function of the number of measurements. The proposed kernel-based approach requires only 60 measurements to reach the same neighborhood of the optimum that the standard approach reaches with 100 measurements.
@@
Figure 3: The optimizer state $\widehat{\theta}$ for both ESC approaches as a function of the number of update steps. The third, fifth, and seventh update in the proposed kernel-based approach are kernel-based update steps, which do not require additional measurements to be performed to perform the update and allow searching for a suitable optimizer gain, resulting in needing only 18 update steps to reach the same neighborhood of the optimum as the standard approach does with 25 update steps.
@@
Figure 4: Contour plot of the steady-state input-output map $f(\theta)$, along with inputs $\widetilde{\theta}_{i}$ applied during measurements in both approaches. These applied inputs lie close to the line passing through the initial optimizer state $\widehat{\theta}_{0}$ and the minimizer $\theta^{*}$ (shown in green).
@@
Figure 5: Cross-section of the input-output map $f(\theta)$ as well as the approximation $m_{k}(\theta)$ at different kernel-based update steps, along the search direction (cf. green line in Figure 4). $m_{k}(\theta)$ quickly improves, allowing large kernel-based update steps to be taken from the circle to the square.
@@ 6 Conclusion
Motivated by the potentially costly and time-consuming nature of performing measurements in practical applications, we presented a novel extremum-seeking approach aimed at reducing the total number of measurements needed to optimize performance. The proposed approach uses kernel-based function approximation to construct online an approximation of the steady-state input-output map of the system, based on data collected during regular extremum-seeking steps. We use this approximation to determine a suitable search direction and optimizer gain to perform a parameter update without performing additional measurements whenever it is sufficiently accurate to guarantee a decrease in the input-output map. If such a decrease cannot be guaranteed, we perform a regular extremum-seeking parameter update step to update the parameters. In addition, we use the data collected during this update step to improve the approximation for the next parameter update step. By using this approach, reductions in both the number of performed measurements and the number of update steps can be obtained, as illustrated by a simulation study in which the number of measurements and updates are reduced by 40% and 28%, respectively.
Future research could consider extensions to a more general class of optimization methods as in Khong et al. (2013a, b), for example, for both the standard parameter update steps and kernel-based update steps. Other potential research directions include the use of variable waiting times instead of a fixed one, as, e.g., in Poveda and Teel (2017), to reduce the time-consuming nature of performing measurements even further, or extensions in which only a subset of all collected data are used, e.g., to reduce computational complexity or to deal with input-output maps that are slowly changing over time.
@@ Appendix A Proof of Theorem 10
We start the proof by showing that under the conditions in Theorem 10, the bound $\bar{b}(\theta,\mu)$ in (37) is indeed the maximum value that $f(\theta^{+})-f(\theta)+c\mu\nabla f(\theta)^{\intercal}\nabla m_{k}(\theta)$ can obtain given the input $\theta\in\mathbb{R}^{n_{\theta}}$. That is, we will start by showing that (37) is the same as
$\bar{b}(\theta,\mu)=$
@@
:=$\mathcal{H}^{\parallel}:=\{f\in\;$ $\mathcal{H}\mid f\in{\rm span}\{\kappa(\cdot,\widetilde{\theta}_{1}),\dots,\kappa(\cdot,\widetilde{\theta}_{N_{k}}),\kappa(\cdot,\theta),$
@@
Let :=$\mathcal{H}^{\perp}:=\{g\in\mathcal{H}\mid\langle g,f^{\parallel}\rangle=0,\,\forall f^{\parallel}\in\mathcal{H}^{\parallel}\}$ be the subspace of functions in $\mathcal{H}$ that are orthogonal to $\mathcal{H}^{\parallel}$ in (48). It then follows that $\mathcal{H}=\mathcal{H}^{\parallel}\oplus\mathcal{H}^{\perp}$, with $\oplus$ denoting the vector direct sum, and that for all $f\in\mathcal{H}$ there exist $f^{\parallel}\in\mathcal{H}^{\parallel}$ and $f^{\perp}\in\mathcal{H}^{\perp}$ such that $f=f^{\parallel}+f^{\perp}$. Using this decomposition of $f$, and observing that $\langle f^{\perp},\kappa(\cdot,z)\rangle$ and $\langle f^{\perp},D^{(0,e_{j})}\kappa(\cdot,\theta)\rangle$ are zero for $z\in\{\widetilde{\theta}_{1},\widetilde{\theta}_{2},\dots,\widetilde{\theta}_{N_{k}},\theta,\theta^{+}\}$ and $j=1,2,\dots,n_{\theta}$, respectively, since for those $z$ and $j$ it holds that $\kappa(\cdot,z)\in\mathcal{H}^{\parallel}$, and $D^{{(0,e_{j})}}\kappa(\cdot,\theta)\in\mathcal{H}^{\parallel}$, we obtain using (49) and (50) that (47a) can be written as
@@
$\>\lvert f^{\parallel}(\widetilde{\theta}_{i})-\widetilde{y}_{i}\rvert\leq\bar{\delta}\;\forall i\in\{1,2,\dots,N_{k}\}.$ (54c)
@@
Step 2: Next, we address the attainment of the supremum by showing that the constraints in (54) define a closed and bounded feasible set. First, note that $\lVert f^{\parallel}\rVert^{2}\leq\Gamma^{2}$ defines a closed and bounded set, since it is the sublevel set of a norm. Next, we note that since sets of the form $\{a\in\mathbb{R}\mid\lvert a\rvert\leq b\}$ are closed in $\mathbb{R}$, so are $\{f^{\parallel}(\widetilde{\theta}_{i})\in\mathbb{R}\mid\lvert f^{\parallel}(\widetilde{\theta}_{i})-\widetilde{y}_{i}\rvert\leq\bar{\delta}\}$ $\forall i\in\{1,2,\dots,N_{k}\}$. Furthermore, the evaluation functional $L_{z}(f^{\parallel})=f^{\parallel}(z)$ is a linear operator and thus the pre-images $L_{z}^{-1}$ of these closed sets are also closed. Consequently, $\{f^{\parallel}\in\mathcal{H}^{\parallel}\mid\lvert f^{\parallel}(\widetilde{\theta}_{i})-\widetilde{y}_{i}\rvert\leq\bar{\delta}\}$ $\forall i\in\{1,2,\dots,N_{k}\}$ are closed in $\mathcal{H}^{\parallel}$. Since the intersection of a finite number of closed sets is necessarily closed, the constraints (54b)-(54c) define a closed feasible set. Moreover, this set is bounded because it is contained in the bounded set defined by the constraint (54b). Since $\mathcal{H}^{\parallel}$ in (48) is finite dimensional, any closed and bounded subset is compact by the Heine-Borel theorem (see, e.g., Royden and Fitzpatrick (2010, Theorem 20)). Therefore, by the extreme value theorem, the continuous objective (54a) attains a maximum on the feasible compact set. Moreover, from Step 1 it follows that the optimizer for which this maximum is obtained must be in $\mathcal{H}^{\parallel}$ ($f^{\perp}=0$), whose members $f^{\parallel}\in\mathcal{H}^{\parallel}$ by (48) have the form
@@
with $K_{\Omega}(\cdot)$ as in (27) for the inputs $\Omega=\{\widetilde{\theta}_{1},\widetilde{\theta}_{2},\dots,\widetilde{\theta}_{N_{k}}\}$, $\nabla K(\theta,\cdot)$ as in (35), and $\beta\in\mathbb{R}^{N_{k}+2+n_{\theta}}$ a weight vector.
@@
$\begin{bmatrix}f^{\parallel}(\widetilde{\theta}_{1})&f^{\parallel}(\widetilde{\theta}_{2})&\cdots&f^{\parallel}(\widetilde{\theta}_{N_{k}})&f^{\parallel}(\theta)&f^{\parallel}(\theta^{+})&\nabla f(\theta)^{\intercal}\end{bmatrix}^{\intercal}$
@@
for all $i\in\{1,2,\dots,N_{k}\}$ with $\xi_{i}\in\mathbb{R}^{N_{k}+2+n_{\theta}}$ a vector with $i$-th element equal to one and other elements equal to zero.
Using (57)-(59), and the observation from Step 2 that the objective (54a) attains a maximum on the feasible compact set, we obtain (37) from (54), completing the proof for $\bar{b}(\theta,\mu)$.
The proof that under the conditions in Theorem 10 $\underaccent{\bar}{b}(\theta)$ in (36) is a lower bound on $\nabla f(\theta)^{\intercal}\nabla m_{k}(\theta)$, i.e., that (36) is the same as
$\underaccent{\bar}{b}(\theta)=$
@@
$\lvert f(\widetilde{\theta}_{i})-\widetilde{y}_{i}\rvert\leq\bar{\delta}\quad\forall i\in\{1,2,\dots,N_{k}\},$ (60c)
@@ Appendix B Proof of Lemma 13
The proof of Lemma 13 is as follows. Since by the conditions of the lemma $\bar{b}(\theta,\mu)\leq 0$, it follows from Theorem 10 that the inequality $f(\theta^{+})\leq f(\theta)-c\mu\nabla f(\theta)^{\intercal}\nabla m_{k}(\theta)$ with $c\in(0,1)$ is satisfied. Therefore, using the definitions of $V(\cdot)$ and $\Delta V(\cdot)$ in (16) and (41), we obtain that the inequality
@@
holds. Furthermore, since by the conditions of the lemma $\underaccent{\bar}{b}(\theta)>0$ for all $\𝒞\theta\in\mathbb{R}\backslash\mathcal{C}$, and by Assumption 2(iii) $\nabla f(\theta)=0$ if and only if $\theta\in\mathcal{C}$, it follows from Theorem 10 and (61) that $\Delta V(\theta)<0$ for all $\theta\in\mathbb{R}^{n_{\theta}}\backslash\mathcal{C}$ and $\Delta V(\theta)=0$ for all $\theta\in\mathcal{C}$. Hence, for any $\Delta_{\theta}$ such that $\lVert\theta\rVert_{\mathcal{C}}\leq\Delta_{\theta}$, there exists a class-$\mathcal{K}$ function $\bar{\rho}_{k}$ such that $\bar{\rho}_{k}(\lVert\theta\rVert_{\mathcal{C}})\leq-\Delta V(\theta)$ (see, e.g., Khalil (2002, Lemma 4.3)). Moreover, it follows from Assumption 14 that $\lVert\theta\rVert_{\mathcal{C}}\geq\omega_{2}^{-1}(V(\theta))$ and thus that $-\bar{\rho}_{k}(\lVert\theta\rVert_{\mathcal{C}})\leq-\bar{\rho}_{k}(\omega_{2}^{-1}(V(\theta)))$. Therefore, we obtain from (61) that during kernel-based extremum-seeking steps the inequality